# On Lan-Sheng-Zuo conjecture

### v2

@@ Acknowledgement
The authors would like to express gratitude to Jianping Wang, who motivated the proof of Lemma 4.3.1 and Proposition 5.1, and to Xucheng Zhang, for discussions about stack. Also, the authors would like to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments that improved the paper.
@@ 5. Lan-Sheng-Zuo conjecture on higher dimensional varieties
In this section we collect some partial results on Lan-Sheng-Zuo conjecture on higher dimensional varieties. Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$.
If $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$, then by Theorem 4.1.1, every $\omega$-semistable vector bundle is strongly semistable. Thus a sufficient method to prove Lan-Sheng-Zuo conjecture on $X$ is to prove every $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ is semistable as a vector bundle.
One method to prove every $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$ is semistable as bundle is to use following estimates:
Lemma 5.1 ([Lan15a, Lemma 5]).
Let $(\mathcal{V},\nabla)$ be a $\nabla$-semistable flat bundle of rank $r$. Then
$L_{\max}(\mathcal{E})-L_{\min}(\mathcal{E})\leq(r-1)\gamma,$
where
$L_{\omega,\max}(\mathcal{E})$ $=\lim_{t\to\infty}\frac{\mu_{\omega,\max}((F^{*})^{k}\mathcal{E})}{p^{k}}$
$L_{\omega,\min}(\mathcal{E})$ $=\lim_{t\to\infty}\frac{\mu_{\omega,\min}((F^{*})^{k}\mathcal{E})}{p^{k}}.$
and $\gamma=(r-1)\max\{L_{\max}(\Omega_{X}^{1}),0\}$.
Lemma 5.2 ([Lan04, Corollary 6.2]).
If $p\geq n=\dim X$ and $L_{\max}(\Omega_{X}^{1})\geq 0$, then
$L_{\max}(\Omega_{X}^{1})\leq\frac{p}{p+1-n}\mu_{\max}(\Omega_{X}^{1}).$
@@
Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$ such that $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$.
  1. $(1)$
    If $\Omega_{X}^{1}$ is strongly $\omega$-semistable, then Lan-Sheng-Zuo holds on $X$.
  2. $(2)$
    If $p\geq n$, then Lan-Sheng-Zuo conjecture holds on $X$.
@@ 2. $(2)$
Firstly, we prove under assumptions $(1)$ and $(2)$, we must have $L_{\omega,\max}(\Omega_{X}^{1})\leq 0$. Indeed, for the case (1), if $\Omega_{X}^{1}$ is strongly $\omega$-semistable, then $L_{\omega}(\mathcal{E})=\mu_{\omega}(\mathcal{E})\leq 0$. For the case (2), if $p\geq n$ and $L_{\omega,\max}(\Omega_{X}^{1})>0$, then by Lemma 5.2 we have
$0<L_{\max}(\Omega_{X}^{1})\leq\frac{p}{p+1-n}\mu_{\max}(\Omega_{X}^{1})\leq 0,$
a contradiction.
If $L_{\omega,\max}(\Omega_{X}^{1})\leq 0$, then for any $\nabla$-semistable flat bundle $(\mathcal{V},\nabla)$, by Lemma 5.1 we have $\mu_{\omega,\max}(\mathcal{V})-\mu_{\omega,\min}(\mathcal{V})=0$ since we always have
$0\leq\mu_{\omega,\max}(\mathcal{V})-\mu_{\omega,\min}(\mathcal{V})\leq L_{\omega,\max}(\mathcal{V})-L_{\omega,\min}(\mathcal{V}).$
Then $\mathcal{V}$ is semistable and thus by Theorem 4.1.1 it is strongly semistable. As a consequence, Lan-Sheng-Zuo conjecture holds on $X$ since we can always choose trivial filtration to run the Higgs-de Rham flow. ∎
For the case (1) in Proposition 5.1, it can be also proved by the same technique used in the second proof of Proposition 3.2.1.
Another proof of (1) in Proposition 5.1.
Let $(\mathcal{V},\nabla)$ be a $\nabla$-semistable flat bundle on $X$. Suppose $\mathcal{V}$ is unstable and its Harder-Narasimhan filtration is given by
@@
where $\mathcal{G}_{i}=\mathcal{V}_{i}/\mathcal{V}_{i-1}$ is $\omega$-semistable bundle for each $i$ with strictly decreasing slopes. Since $\mathcal{V}$ is $\omega$-semistable if $(F^{*})^{N}\mathcal{V}$ is $\omega$-semistable, so up to finite times Frobenius pullback, we may assume $\mathcal{G}_{i}$ is strongly $\omega$-semistable for each $i$.
Since $(\mathcal{V},\nabla)$ is $\nabla$-semistable, we must have the maximally destabilizer $\mathcal{V}_{1}$ is not $\nabla$-invariant. Choose $\ell>1$ such that $\nabla\left(\mathcal{V}_{1}\right)\subseteq\mathcal{V}_{\ell}$ but $\nabla\left(\mathcal{V}_{1}\right)\nsubseteq\mathcal{V}_{\ell-1}$. Then $\nabla$ induces a non-zero morphism $\mathcal{G}_{1}\to\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$. But if $\ell\neq 1$, there is no non-zero morphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ since $\mu_{H}(\Omega_{X}^{1})\leq 0$ and both $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ are semistable, as tensor product of strongly semistable vector bundle is again strongly semistable. This shows $\mathcal{V}_{1}$ is $\nabla$-invariant, a contradiction. ∎
Theorem 5.1 ([Lan15b, Theorem 0.1]).
Let $X$ be a smooth $n$-dimensional projective variety over $k$ with $\text{char}(k)=p\geq(n-1)(n-2)$. Assume $K_{X}$ is numerically trivial and $X$ is not uniruled. Then for every ample divisor $\omega$ the tangent bundle of $X$ is strongly $\omega$-semistable and $c_{2}(X)\cdot\omega^{n-2}\geq 0$.
Theorem 5.2 ([Lan15b, Theorem 4.5]).
Let $X$ be a K3 surface over $k$. Then the tangent bundle of $X$ is strongly semistable with respect to some non-trivial nef polarization.
Corollary 5.1.
Lan-Sheng-Zuo conjecture holds on
  1. $(1)$
    abelian varieties;
  2. $(2)$
    K3 surface with respect to some non-trivial nef polarization;
  3. $(3)$
    smooth projective Calabi-Yau varieties over $k$ with $\text{char}(k)=p\geq(n-1)(n-2)$, which is not uniruled, with respect to every ample polarization.
Proof.
For $(1),(2)$ and $(3)$, all of them have strongly $\omega$-semistable tangent bundle, and thus by Proposition 5.1 the Lan-Sheng-Zuo conjecture holds. ∎
Remark 5.1.
Besides cases in Proposition 5.1 and counter-example on certain smooth projective of general type, the Lan-Sheng-Zuo conjecture on higher dimensional varieties is still unknown. By results for curve case, we shall expect the Lan-Sheng-Zuo conjecture holds for all smooth projective varieties which are Fano or Calabi-Yau, and fails for all smooth projective varieties which are general type.

### v3

@@ Acknowledgement
The authors would like to express gratitude to Jianping Wang, who motivated the proof of Lemma 4.3.1, and his comments on higher dimensional case, which improve the Proposition 5.1. Also, the authors would like to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments that improved the paper.
@@ 5. Lan-Sheng-Zuo conjecture on higher dimensional varieties
In this section we prove a case of the Lan-Sheng-Zuo conjecture on higher dimensional varieties.
@@
Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$ such that $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$. Then the Lan-Sheng-Zuo conjecture holds on $X$.
@@
Let $(\mathcal{E},\theta)$ be a semistable Higgs bundle of exponent $\leq p-1$ and $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$. Then $(\mathcal{V},\nabla)$ is a $\nabla$-semistable flat bundle. Now it suffices to prove that $\mathcal{V}$ is semistable as a vector bundle. If so, then by Theorem 4.1.1 we have $\mathcal{V}$ is strongly semistable, and thus we can always choose trivial filtration to run the Higgs-de Rham flow.
Suppose $\mathcal{V}$ is unstable and its Harder-Narasimhan filtration is given by
@@
where $\mathcal{G}_{i}=\mathcal{V}_{i}/\mathcal{V}_{i-1}$ is $\omega$-semistable bundle for each $i$ with strictly decreasing slopes. Since $(\mathcal{V},\nabla)$ is $\nabla$-semistable, we must have the maximally destabilizer $\mathcal{V}_{1}$ is not $\nabla$-invariant. Choose $\ell\in\operatorname{\mathbb{Z}}_{>1}$ such that $\nabla\left(\mathcal{V}_{1}\right)\subseteq\mathcal{V}_{\ell}$ but $\nabla\left(\mathcal{V}_{1}\right)\nsubseteq\mathcal{V}_{\ell-1}$. Then $\nabla$ induces a non-zero morphism $\mathcal{G}_{1}\to\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$.
Suppose the maximal destabilizer of $\Omega_{X}^{1}$ is $(\Omega_{X}^{1})_{1}$. Then by Theorem 4.1.1 we have both $(\Omega_{X}^{1})_{1}$ and $\mathcal{G}_{i}$ are strongly semistable, and thus the tensor product $\mathcal{G}_{\ell}\otimes(\Omega_{X}^{1})_{1}$ is semistable. As a consequence, we have
$\mu_{\omega,\max}(\mathcal{G}_{\ell}\otimes\Omega_{X}^{1})$ $\leq\mu_{\omega}(\mathcal{G}_{\ell})+\mu_{\omega,\max}(\Omega_{X}^{1})$
$\leq\mu_{\omega}(\mathcal{G}_{\ell})$
$<\mu_{\omega}(\mathcal{G}_{1}).$
This shows there is no non-zero morphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ for $\ell\in\operatorname{\mathbb{Z}}_{>1}$, a contradiction. ∎